X-N(45, variance) and p(x>51)=0.288 find standard deviation

1
Expert's answer
2022-05-02T15:37:08-0400
"P(X>51)=1-P(Z\le\dfrac{51-45}{\sigma})=0.288"

"\dfrac{51-45}{\sigma}=0.559237"

"\sigma=\dfrac{6}{0.559237}\approx10.7289"